solve by the elimination method
0.5x+0.25y=66
0.15x+0.05y=72
Ijust do not understand this could someone help please? Thanks

Answer by checkley77(12569)

0.5x+0.25y=66 multiply by -.3 & add.
0.15x+0.05y=72
-0.15x-0.075y=-19.8
0.15x+.05y=72
-----------------------
-.025y=52.2
y=52.2/-.025
y=-2088 ans.
.5x+.25*-2088=66
.5x-522=66
.5x=66+522
.5x=588
x=588/.5
x=1,176 ans.